simplify:
(square root of 7) x (square root of 14)

Answer by checkley75(3666)

SQRT7*SQRT14=SQRT7*14=SQRT98=9.8994949 ANSWER.
OR SQRT7*SQRT(2*7)=7SQRT2=7*1.4142135=8.8994949 ANSWER.